14. a) RBC stock has a beta of 0.85, while TD stock has a beta of 1.21. The risk-free rate is 1.5%, and the expected return on a portfolio with 50% weight in RBC and the remainder in TD is 9.74%. What is the market risk premium?

b) You are offered an opportunity to invest in a stock that has a beta of 1.5. If the risk-free rate is 2.4%, and the expected market return is 10%, what is the minimum return that you should accept for doing so?
Answer #1

1.remainder investment in TD=100-50=50%

Expected rate=risk free rate+Beta*market risk premium

Expected rate for:

RBC=1.5+(0.85*market risk premium)

TD=1.5+(1.21*market risk premium)

Portfolio return=Respective return*Respective weight

9.74=0.5*[1.5+0.85*market risk premium]+0.5[1.5+1.21*market risk premium]

9.74=0.75+0.425*market risk premium+0.75+0.605*market risk premium

market risk premium=(9.74-0.75-0.75)/(0.425+0.605)

=8%

2.Required return=risk free rate+beta*(market rate-risk free rate)

=2.4+1.5*(10-2.4)

=13.8%